OpenStudy (anonymous):

find the inverse laplace of 2s exp^-s/(s-1)^2

OpenStudy (luigi0210):

Could you rewrite that using the equation tool?

OpenStudy (anonymous):

y u dont understand?

OpenStudy (anonymous):

\[(2s \times e^{-s})/(s-1)^2\] ?

OpenStudy (anonymous):

yes this is right.

OpenStudy (luigi0210):

Have you tried anything so far?

OpenStudy (anonymous):

yes but im little bit confuse.

OpenStudy (luigi0210):

Show us what you've tried so far and we'll see if we can help

OpenStudy (anonymous):

ok my ans is 2exp^t-1.H(t-1) but the book ans is different.

OpenStudy (luigi0210):

What's the book answer?

OpenStudy (anonymous):

2t.exp^t-1H(t-1).

OpenStudy (luigi0210):

Did you do this with partial fractions?

OpenStudy (anonymous):

no.i use unit step function.

hartnn (hartnn):

After applying time shifting, you'll need to take inverse laplace of 2s/(s-1)^2 how did you find it? you say you got 2e^(t) for this?

hartnn (hartnn):

lets see :) Applying frequency shifting, \(\Large L^{-1}[ 2s/ (s-1)^2] = e^t L^{-1} [2(s+1)/s^2 ]\) clear till here ?

hartnn (hartnn):

because next steps are quite trivial !

OpenStudy (anonymous):

the book ans is 2t.exp^t-1H(t-1).

OpenStudy (anonymous):

i need this answer.

hartnn (hartnn):

lets see how we can get there :) you understood the properties i applied? Have the inverse laplace transform table with you ?

OpenStudy (anonymous):

yes i have.

hartnn (hartnn):

so from 2s exp^-s/(s-1)^2 we came to 2 (s+1)/s^2 by applying 2 properties. you understood both ?

hartnn (hartnn):

for better clarity, do you want me to give you the steps till that step?

OpenStudy (anonymous):

yes give me details plz

hartnn (hartnn):

give me a moment to type, by that time, please think of next steps :)

OpenStudy (anonymous):

ok.thank u.

hartnn (hartnn):

Since \(\Large L^{-1} [e^{-as}F(s)] = f(t-a) H(t-a) \) we first find \(\Large L^{-1} [\dfrac{2s }{(s-1)^2}] \) and then use the above property Now \(\Large L^{-1}f(s-a) = e^{at}f(t) \) so \(\Large L^{-1} [\dfrac{2s }{(s-1)^2}] = e^t L^{-1} [\dfrac{2(s+1) }{(s)^2}] \) see whether you get these steps first :)

hartnn (hartnn):

\(\Large L^{-1}[f(s-a)] = e^{at}L^{-1} [f(s)]\)

OpenStudy (anonymous):

ok thank u so much.and which softwaer u used to write this things clearly.

OpenStudy (anonymous):

last thing is tell me the the laplace of exp^t-1 is what??

hartnn (hartnn):

used \(Latex \), which is supported by this site. you can use it too :)

OpenStudy (anonymous):

can i download it?/

hartnn (hartnn):

why would you need that ? we're here : \(e^t L^{-1} [\dfrac{2(s+1) }{(s)^2}]\) you'll get the same answer which is in you book, be patient lol we're just half way done

hartnn (hartnn):

and for the latex you can just start typing in `\( ..\)` try it type `\(Hi\)` and see what u get

OpenStudy (anonymous):

ok

hartnn (hartnn):

so we need to solve \(e^t L^{-1} [\dfrac{2(s+1) }{(s)^2}]\) `\(e^t L^{-1} [\dfrac{2(s+1) }{(s)^2}]\)` any ideas how to proceed?

OpenStudy (anonymous):

sorry exp ^t-x laplace??

hartnn (hartnn):

sorry, didn't get your question. we're solving an inverse laplace question, and I just asked you whether you know how we can proceed with \( L^{-1} [\dfrac{2(s+1) }{(s)^2}]\) are we on a same page?

OpenStudy (anonymous):

i dont understand why u right 2(s+1) in numerater?

hartnn (hartnn):

glad that you ask doubts :) \(\Large L^{-1}[f(s-a)] = e^{at}L^{-1} [f(s)]\) This is a standard property what does it say? If you want to remove e^at from inverse laplace , replace 's' by 's+1' in your function <<Note how s-1 on left became s>>

hartnn (hartnn):

so applying that property to your function, the numerator s became (s+1)

hartnn (hartnn):

correction : If you want to remove e^at from inverse laplace , replace 's' by 's+a' in your function <<Note how s-a on left became s>>

OpenStudy (anonymous):

can u right the correct ans which i post in above comments??

hartnn (hartnn):

if I just post all the steps, you'll understand none of them. I don't suppose you want that... thats why I am posting few steps and asking you whether you got them... if this approach doesn't suit you, I am really not the right person to help you...

OpenStudy (anonymous):

ok thank u so much to give me precious time.nice to meet u.

hartnn (hartnn):

ah, you too :)

OpenStudy (sidsiddhartha):

here \[F(s)=\frac{ 2s.e^{-s} }{ (s-1)^2 }\\s~In ~laplace~domain~means~\frac{ d }{ dt }~In ~time~domain\\f(t)=2.\frac{ d }{ dt }L^{-1}[\frac{ e^{-s} }{ (s-1)^2 }]=2.\frac{ d }{ dt }[(t-1)e^{t-1}H(t-1)]\\used~frequency~shifting~In~last~step\\f(t)=2[(t-1)e^{t-1}+e^{t-1}]H(t-1)=2t.e^{t-1}H(t-1)\] easy this way :) @Yasirist